OpenStudy (anonymous):

FAN AND MEDAL FOR HELP! Given the equation |2x+3| = x-5 Which of the followiung would be the solution set for x? A. {-8} B. {2/3} C. {-8, 2/3} D No Solution E None of These @amistre64

OpenStudy (amistre64):

test them out

OpenStudy (anonymous):

I got for my answers |-13| = -13 and |13/3| = -13/3 So would the answer be C or No solution?

OpenStudy (amistre64):

does |-a| = -a ?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Isn' abs value supposed to be equal to a positive? Im rusty lol

OpenStudy (amistre64):

absolute values are never negative ... so x < 5 is not going to produce any valid results. so, either a solution exists; or none exists ... which do you think it should be and why?

OpenStudy (anonymous):

I'm assuming none since |-a| = a and \[a \neq a\]

OpenStudy (anonymous):

a =/ -a **

OpenStudy (amistre64):

none exist ... either: 2x+3 = x-5 , such that x>5 x=-8 or -(2x+3) = x-5, such that x > 5 2 = 3x x=2/3 neither 'solution' is valid

OpenStudy (amistre64):



OpenStudy (anonymous):

Thanks!